OpenStudy (anonymous):

Given f(x)=3-x and g(x)=x^2-x, what is f(g(x)? Can someone show me too?

OpenStudy (amistre64):

just fit g(x) into the 'x' for f(x) like this: f(g(x)) = 3 - g(x)

OpenStudy (amistre64):

f(g(x)) = 3 - (x^2 - x) f(g(x)) = 3 - x^2 + x)

OpenStudy (anonymous):

Oooh okay! I thought it was something like that.

OpenStudy (anonymous):

I just forgot to put it on the other side.

OpenStudy (anonymous):

Thank you.